OpenStudy (anonymous):

A) Write an equation of a line that passes through the point (2, 4) and whose slope is 1/2. B)Find the equation of a line in slope-intercept from that passes through the points (6,5) and (4,1). C)Write the equation of the horizontal or vertical line: (-7, 3), slope is undefined. D)Find the equation of a line perpendicular to the given line: y = 3x – 1. E)Write an equation in slope-intercept form for an equation of a line perpendicular to y = -3x +1 and that intersects at point (1, 2/3).

OpenStudy (anonymous):

a.)use: \[y=mx+b\],where m is the slope, solve for the value of b using the given coordinates b.)use the slope formula\[m=\frac{ \Delta y }{ \Delta x }\], then solve for b again

OpenStudy (anonymous):

A.) y - y1 = m(x - x1) using slope 1/2 and points (2,4) y - 4 = 1/2(x - 2) y - 4 = 1/2x - 1 y = 1/2x - 1 + 4 y = 1/2x + 3 B) find the slope first slope(m) = (y2 - y1) / (x2 - x1) using points (6,5) and (4,1) m = (1 - 5) / (4 - 6) m = -4/-2 m = 2 y - y1 = m(x - x1) using slope 2 and points (6,5) y - 5 = 2(x - 6) y - 5 = 2x - 12 y = 2x - 12 + 5 y = 2x - 7 C) It is a vertical line. The equation is y = 3 D) A perpendicular line will have a negative reciprocal of the slope given. The slope given is 3, so the slope you need will be -1/3. I am sorry I am not sure about this one ....because there is no points given. E) perpendicular line...negative reciprocal to slope -3 is 1/3. y - y1 = m(x - x1) y - 2/3 = 1/3(x - 1) y - 2/3 = 1/3x - 1/3 y = 1/3x - 1/3 + 2/3 y = 1/3x + 1/3 any questions ?